If the polynomial x2-15-m(2x-8) has equal zeroes find the value of m

Grade:10

1 Answers

Nilkanth Sawant
7 years ago
x2-2mx+8m-15
x2+(8-2m)x-15
zeroes are equal,therefore b2-4ac=0
(8-2m)2-4 x-15=0
64-32m+4m2+60=0
4m2-32m+124=0
m2-8m+31=0
Solve the equation and get the answer!!!!